# floZ: Evidence estimation from posterior samples with normalizing flows

### v1

@@ IV Conclusions
In this work, we have introduced floZ, a novel method using normalizing flows for estimating the Bayesian evidence from posterior samples. Our approach computes the Bayesian evidence and its numerical uncertainty using pre-existing samples drawn from the target unnormalized posterior distribution, e.g. as obtained by standard MCMC methods.
@@
Normalizing flows are built to map a complex target probability function (in our case, the unnormalized posterior distribution, given by the product of a likelihood and prior) into a simple probability distribution (e.g. a Gaussian). The transformation between the original variables and the latent ones is bijective and can be modeled with a neural network. Importantly for our method, the flows encode the volume of the real-space distribution, i.e. the Bayesian evidence. Our method hinges on this fact to compute the evidence from a set of posterior samples previously gathered by any suitable algorithm (e.g., MCMC), while at the same time minimizing the scatter of the estimates of the evidence across the samples by adding suitable terms to the custom loss function.
@@
We have validated floZ against analytical benchmarks in parameter spaces of up to 15 dimensions: multivariate Gaussians, a finite multivariate Gaussian mixture, as well as the Rosenbrock distribution. Our method demonstrates performance comparable to nested sampling (which requires to be run ad hoc) and superior to a $k$-nearest neighbors method employed on the same posterior samples. We believe that floZ will add to the toolbox of astronomers and cosmologists seeking a fast, reliable method to compute the evidence from existing sets of posterior samples. floZ will be especially useful for cases where the likelihood is expensive and a full nested sampling run is difficult to achieve, while MCMC samples can be obtained more efficiently thanks to the ease of parallelization of many MCMC algorithms (such as Gibbs sampling). Pushing our approach to much larger parameter spaces is challenged by two limitations, both fundamentally stemming from the curse of dimensionality: first, obtaining posterior samples in high dimensions becomes typically more challenging (although some methods, like Gibbs sampling and Hamiltonian Monte Carlo, can show mild scaling with dimensionality in some favorable circumstances); second, training the flow in higher dimensions suffers from increasing inaccuracies. However, we note that we only need accurate densities in latent space near the peak (as opposed to into the tails of the distribution), which opens the door to potentially exploiting the existence of a well-defined ‘typical set’ in high dimensions, a consequence of the concentration of measure phenomenon. We shall explore this idea further in a dedicated future paper.
@@
Among the applications of floZ to gravitational-wave astronomy, it is worth mentioning the recent evidence for a stochastic signal reported by pulsar timing array experiments Antoniadis:2023ott ; Tarafdar:2022toa ; NANOGrav:2023gor ; Reardon:2023gzh ; Xu:2023wog . Within a Bayesian framework, interpreting the nature of this stochastic signal requires comparing different hypotheses and their evidences, which floZ could readily compute from the existing samples released by the experiments. Similarly, as another of the many possible applications, floZ could be used to investigate the statistical robustness of features/peaks that may be present in the mass function of the astrophysical black holes detected by the LIGO-Virgo-KAGRA collaboration KAGRA:2021duu .
@@ Acknowledgements.
We thank Alan Heavens for insightful discussions. E.B., R.S. and M.B. acknowledge support from the European Union’s H2020 ERC Consolidator Grant “GRavity from Astrophysical to Microscopic Scales” (Grant No. GRAMS-815673), the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”, and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855. M.C. is funded by the European Union under the Horizon Europe’s Marie Sklodowska-Curie project 101065440. R.T. acknowledges co-funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 – Project FAIR “Future Artificial Intelligence Research”. This resource was co-financed by the Next Generation EU [DM 1555 del 11.10.22]. RT is partially supported by the Fondazione ICSC, Spoke 3 “Astrophysics and Cosmos Observations”, Piano Nazionale di Ripresa e Resilienza Project ID CN00000013 “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di “campioni nazionali di R&S (M4C2-19 )” - Next Generation EU (NGEU).

### v2

@@ V Conclusions
We have introduced floZ, an improved method that uses normalizing flows for estimating the Bayesian evidence and its numerical uncertainty using pre-existing samples drawn from the target unnormalized posterior distribution, e.g. as obtained by standard MCMC methods.
@@
Normalizing flows are built to map a complex target probability function (in our case, the unnormalized posterior distribution, given by the product of a likelihood and prior) into a simple probability distribution (e.g. a Gaussian). The transformation between the original variables and the latent ones is bijective and can be modeled with a neural network. We exploit the fact that the flows encode the volume of the real-space distribution to compute the evidence from a set of posterior samples previously gathered by any suitable algorithm (e.g., MCMC), while at the same time reducing the estimate’s scatter thanks to three novel additional loss terms for the training of the normalizing flow. Furthermore, we address the limitation of modeling distribution with sharp boundaries with a pre-processing technique showing marked improvement in the results from normalizing flows and other state-of-the-art evidence estimation techniques that use nested sampling and $k$-nearest-neighbours.
@@
We have validated floZ against analytical benchmarks in parameter spaces of up to 15 dimensions: multivariate Gaussians, a finite multivariate Gaussian mixture, multivariate exponential distribution, as well as, the Rosenbrock distribution. Our method demonstrates performance comparable to nested sampling (which requires to be run ad hoc) and superior to a $k$-nearest neighbors method employed on the same posterior samples. We believe that floZ will add to the toolbox of astronomers and cosmologists seeking a fast, reliable method to compute the evidence from existing sets of posterior samples. floZ will be especially useful for cases where the likelihood is expensive and a full nested sampling run is difficult to achieve, while MCMC samples can be obtained more efficiently thanks to the ease of parallelization of many MCMC algorithms (such as Gibbs sampling).
Pushing our approach to much larger parameter spaces remains challenging because of two limitations, both fundamentally stemming from the curse of dimensionality: first, obtaining posterior samples in high dimensions becomes more difficult (although some methods, like Gibbs sampling and HMC, can show mild scaling with dimensionality in some favorable circumstances); second, training the flow in higher dimensions suffers from increasing inaccuracies (but see Piras2024TheFO for a normalizing-flows-enhanced harmonic mean estimator with similar performance as our own up to 150 dimensions, at least for the restricted case of unimodal, mildly non-Gaussian targets). However, we note that we only need accurate densities in latent space near the peak (as opposed to into the tails of the distribution), which opens the door to potentially exploiting the existence of a well-defined ‘typical set’ in high dimensions, a consequence of the concentration of measure phenomenon. We shall explore this idea further in a dedicated future paper.
@@
To demonstrate the use of $floZ$ on real data, we estimated the Bayes factor of the presence/absence of the first overtone in the ringdown phase of the first GW observation by the LIGO-Virgo collaboration GW150914 , namely GW150914. Our estimate is consistent with that of nested sampling. Among other applications of floZ to gravitational-wave astronomy, it is worth mentioning the recent evidence for a stochastic signal reported by pulsar timing array experiments Antoniadis:2023ott ; Tarafdar:2022toa ; NANOGrav:2023gor ; Reardon:2023gzh ; Xu:2023wog . Within a Bayesian framework, interpreting the nature of this stochastic signal requires comparing different hypotheses and their evidences, which floZ could readily compute from the existing samples released by the experiments. Similarly, as another of the many possible applications, floZ could be used to investigate the statistical robustness of features/peaks that may be present in the mass function of the astrophysical black holes detected by the LIGO-Virgo-KAGRA collaboration KAGRA:2021duu .
@@ Acknowledgements.
We thank Alan Heavens for insightful discussions and Uroš Seljak for comments on an earlier draft. E.B., R.S. and M.B. acknowledge support from the European Union’s H2020 ERC Consolidator Grant “GRavity from Astrophysical to Microscopic Scales” (Grant No. GRAMS-815673), the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”, and the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855. M.C. is funded by the European Union under the Horizon Europe’s Marie Sklodowska-Curie project 101065440. R.T. acknowledges co-funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 – Project FAIR “Future Artificial Intelligence Research”. This resource was co-financed by the Next Generation EU [DM 1555 del 11.10.22]. RT is partially supported by the Fondazione ICSC, Spoke 3 “Astrophysics and Cosmos Observations”, Piano Nazionale di Ripresa e Resilienza Project ID CN00000013 “Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di “campioni nazionali di R&S (M4C2-19 )” - Next Generation EU (NGEU).